OpenStudy (anonymous):

Help Does anyone know how to do this. Write 4152 in its expanded form.

OpenStudy (john_es):

May be, \[4\cdot 1000+1\cdot 100+5\cdot 10+2\]

OpenStudy (dls):

^^

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

the answer has to be like this. https://angel.spcollege.edu/AngelUploads/QuestionData/242b909d-e3ac-49a3-88f1-8f0eb606deaa/46534223453245411922.png# {45f01d4a-5edb-40c2-b109-14aefb792253}

OpenStudy (john_es):

May be this helps you, http://www.aaamath.com/g31d_px1.htm

OpenStudy (anonymous):

not sure how to write it with the parentheses

OpenStudy (mathstudent55):

\(4152\) \(= 4\times 1000+1\times 100+5\times 10+2 \times 1 \) Remember that \(1 = 10^0\) \(10 = 10^1\) \(100 = 10^2\) \(1000 = 10^3\) etc. \(= 4\times 1000+1\times 100+5\times 10+2 \times 1 \) \(= 4\times 10^3+1\times 10^2+5\times 10^1+2 \times 10^0 \) \(= (4\times 10^3) + (1\times 10^2) + (5\times 10^1) + (2 \times 10^0) \)

OpenStudy (anonymous):

your the best thank u

OpenStudy (mathstudent55):

You're welcome.